Convert the following frequencies to wavelengths (please present the final answer in the most appropriate units to present the quantities in a succinct manner):

a. 10.37 Hz

b. 400.1 MHz

c. 1.91 GHz

d. 79 GHz

e. 167 MHz


1
Expert's answer
2021-10-15T10:24:52-0400

We know from the start that "c=3 \times 10^8\,m/s=\lambda \cdot \nu \implies \lambda=\cfrac{c}{\nu}". Then we start to calculate the values as it follows


a. 10.37 Hz "\implies \lambda_a=\cfrac{3 \times 10^8\,m/s}{10.37\,Hz}\cdot \cfrac{1\,Hz}{1\,s^{-1}}=2.893\times 10^{7}\,m=289.3\,nm"

b. 400.1 MHz "\implies \lambda_b=\cfrac{3 \times 10^8\,m/s}{400.1\,MHz}\cdot \cfrac{1\,MHz}{10^6\,s^{-1}}=0.750\,m"

c. 1.91 GHz "\implies \lambda_c=\cfrac{3 \times 10^8\,m/s}{1.91\,GHz}\cdot \cfrac{1\,GHz}{10^9\,s^{-1}}=0.157\,m"

d. 79 GHz "\implies \lambda_d=\cfrac{3 \times 10^8\,m/s}{79\,GHz}\cdot \cfrac{1\,GHz}{10^9\,s^{-1}}=3.797\times 10^{-3}\,m=3.797\, mm"

e. 167 MHz "\implies \lambda_e=\cfrac{3 \times 10^8\,m/s}{167\,MHz}\cdot \cfrac{1\,MHz}{10^6\,s^{-1}}=1.796\,m"


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.